OpenStudy (tatumlee):

How many solutions does this system have? x-2y=2 y=-2x+5

Directrix (directrix):

x-2y=2 y=-2x+5 In the top equation, substitute (-2x +5) for y x-2y=2 x - 2*(-2x + 5) = 2 x +4x -10 = 2 5x - 10 = 2 x = ? @tatumlee

OpenStudy (tatumlee):

I added 10 to both sides and ended up with 5x=12 am I doing this right?

OpenStudy (tatumlee):

@Directrix

Directrix (directrix):

>> 5x=12 am I doing this right? Yes. x = 12/5 That indicates that there is a solution to the system of equations.

Directrix (directrix):

y=-2x+5 y = (-2)* 12/5 + 5 y = 1/5 The single solution is (12/5 , 1/5)

Directrix (directrix):

@tatumlee